OpenStudy (anonymous):

can anyone help me through with the steps to solve for w?? 8-w=12

OpenStudy (anonymous):

W=-4

OpenStudy (anonymous):

Think, you cant subtract a positive and get a bigger number, so it would have to be a negative. at that point, you would just need to look at the difference

OpenStudy (anonymous):

Let's solve your equation step-by-step. 8−w=12 Step 1: Simplify both sides of the equation. 8−w=12 Simplify: −w+8=12 Step 2: Subtract 8 from both sides. −w+8−8=12−8 −w=4 Step 3: Divide both sides by -1. −w/−1=4/−1 w=−4 Answer: w=−4

OpenStudy (anonymous):

8-w=12 8-12=w -4=w so, w=-4